I have having great difficuly with these worded problems, your help would be greatly appreciated. Thank you.
It asks me:
To solve the following word problem using quadratic equation:
Then goes on to state:
Find the dimensions of a rectangular field that has a perimeter of 724m and an area of 32400m^2.
Thank you.
Cheers D.

Answer by stanbon(75887)

Find the dimensions of a rectangular field that has a perimeter of 724m and an area of 32400m^2.
-------------------------------------
Perimeter = 2(length + width)
724 = 2(length + width)
length + width = 362
--------------------------
length * width = 32400
================================
EQUATIONS:
l = 362-w
Substitute to solve for width:
(362 -w)w = 32400
-w^2+362w-32400
w^2-362w+32400 = 0
w = [362 +- sqrt(1444)]/2
w = [362 +- 38]/2
w = 200 (width)
---------------
Solve for length:
l + 200 = 362
l = 162 (length)
=================================
Cheers,
Stan H.